40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 28 13:32:01 2010, alberto, kiwamu, Update, VAC, vacuum work 
    Reply  Fri May 28 13:51:51 2010, Jenne, Update, VAC, vacuum work 
       Reply  Fri May 28 14:24:30 2010, Jenne, Update, VAC, vacuum work 
Message ID: 3009     Entry time: Fri May 28 13:32:01 2010     Reply to this: 3010
Author: alberto, kiwamu 
Type: Update 
Category: VAC 
Subject: vacuum work 

We started a vacuum work in this morning. And still it's going on.

 

Although the last night the green team replaced a steering mirror by an 80% reflector on the PLS table, the beam axis to the MC looks fine.

The MC refl beam successfully goes into the MCrefl PD, and we can see the MC flashing as usual.

We started measuring the distance of the optics inside the vacuum chamber, found the distance from MC3 to MMT1(curved mirror) is ~13cm shorter than the design.

We moved the positions of the flat mirror after the Faraday and the MMT1, but could not track the beam very well because we did not completely lock the MC.

Now we are trying to get the lock of the MC by steering the MC mirrors.

 

P.S.

Kevin suceeded in locking it !!

ELOG V3.1.3-